two particles A&B are in motion.if wavelength associated with particle A is 5*10(power=-8)m.calculate wavelength associated with particle B if its momentum is half of A....

Grade:11

1 Answers

Ashutosh Sharma
8 years ago
(mV)A=h/λ(A)---- (1)
(mV)B=h/λ(B)------- (2)
so from 1&2 (mV)B=12(mV)A
therefore,
hλB=h2λA=2×5×10−8m=10×10−8m=1000A∘